# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 2.1. The Bruhat-Tits building and Weyl groups
We let $\mathscr{X}$ denote the semisimple Bruhat-Tits building of $G$. In the standard apartment $X_{*}(\mathbb{T}/\mathbb{C})\otimes_{{\mathbb{Z}}}\mathbb{R}$ of $\mathscr{X}$, we fix a chamber $C$ and a hyperspecial vertex $x_{0}$ such that $x_{0}\in\overline{C}$. Given a face $F$ in $\mathscr{X}$, we let $P_{F}$ denote the parahoric subgroup of $G$ associated to $F$ and $P_{F}^{\dagger}$ be the stabiliser in $G$ of the face $F$. The group $P_{F}$ has a pro-$p$ radical which we denote by $I_{F}$. In particular, we set $I=I_{C}$ and $I^{\prime}=P_{C}$ which are called a pro-$p$ Iwahori subgroup and Iwahori subgroup of $G$, respectively.
The chamber $C$ determines a set $\Phi^{+}$ of positive roots of the root system $\Phi\subseteq X^{*}(\mathbb{T}/\mathbb{C})$ of $(\mathbb{G},\mathbb{T})$. We will view elements of $\Phi$ as characters of $T$. If $\Phi^{\prime}$ is an irreducible component of $\Phi$, we write $\mathrm{rk}(\Phi^{\prime})$ to denote its rank. We let $\Pi$ denote the basis of $\Phi$ defined by $\Phi^{+}$, and $B=T\ltimes U$ be the Borel subgroup containing $T$ defined by $\Phi^{+}$, where $U$ is the unipotent radical of $B$. A standard parabolic subgroup $P=M\ltimes N$ is any parabolic subgroup of $G$ containing $B$. The corresponding Levi subgroup $M$ will also be called standard. Given a standard parabolic subgroup $P=MN$, we let $\Pi_{M}$ (resp. $\Phi_{M}$, resp. $\Phi_{M}^{+}$) be the corresponding simple roots (resp. root system, resp. positive roots) defined by $M$. We will sometimes also write $\Pi_{P}$ to denote this same set of simple roots.
@@
the extended Weyl group of $(G,T)$, where $\Lambda:=T/(T\cap I^{\prime})$. The action of $G$ on $\mathscr{X}$ restricts to an action of $W$ on the standard apartment by affine automorphisms. We let
@@
Let $\Omega=\{w\in W\mid\ell(w)=0\}$. Then we have a decomposition $W=W_{\text{aff}}\rtimes\Omega$ where $W_{\text{aff}}$ denotes the affine Weyl group, generated by the set $S$ of simple affine reflections fixing the walls of $C$. Moreover, if we let $\Pi_{\operatorname{aff}}=\Pi\cup\{(\alpha_{1},1),\ldots,(\alpha_{r},1)\}\subseteq\Phi^{+}_{\operatorname{aff}}$ where $\alpha_{1},\ldots,\alpha_{r}$ range over the longest roots in each irreducible component of the root system $\Phi$, then the elements of $S$ are all of the form $s_{\alpha}$ for $\alpha\in\Pi_{\operatorname{aff}}$ (c.f. [25, §4.3]). Given any affine root $\alpha$ of the form $\alpha(t)=\beta(t)+\lambda$ with $\beta\in\Phi$, $\lambda\in{\mathbb{Z}}$, and $t\in T$, we set $\alpha^{\vee}=\beta^{\vee}$. For any given $s=s_{\alpha}\in S$, we let the subtorus $T_{s}\subseteq T$ denote the $\mathfrak{F}$-rational points of the algebraic group image of the cocharacter $\alpha^{\vee}$. Its group of $\mathbb{F}_{q}$-rational points $T_{s}(\mathbb{F}_{q})$ is a subgroup of $T(\mathbb{F}_{q})$ isomorphic to $\mathbb{F}_{q}^{\times}$. Now, given an $\overline{\mathbb{F}_{p}}$-valued character $\xi$ of $T(\mathbb{F}_{q})$, we let
(2.2) $S_{\xi}:=\{s\in S\mid\xi(t)=1\text{ for all }t\in T_{s}(\mathbb{F}_{q})\}.$
@@
To each face $F\subseteq\overline{C}$ we associate the subset $S_{F}\subseteq S$ of those reflections fixing $F$ pointwise and we let $W_{F}$ denote the subgroup of $W$ generated by $S_{F}$. Note that $S_{C}=\emptyset$. If $\Phi=\sqcup_{i=0}^{r}\Phi_{i}$ is the decomposition of $\Phi$ into irreducible components and $\Pi=\sqcup_{i=0}^{r}\Pi_{i}$ is the corresponding decomposition of the basis, we also have that $S=\sqcup_{i=0}^{r}S_{i}$ where, in the above notation, $S_{i}=\{s_{\beta}\mid\beta\in\Pi_{i}\}\cup\{s_{(\alpha_{i},1)}\}$. Then the map
@@
is a bijection (c.f. [8, §1.3.5]). Moreover, this bijection is order-reversing in the sense that $F^{\prime}\subseteq\overline{F}$ if and only if $S_{F}\subseteq S_{F^{\prime}}$.
@@ 2.2. Pro-$p$ Iwahori-Hecke algebras
where $C(F)$ is the chamber defined in [18, Lemma 1.3], and the corresponding Hecke algebras are then defined analogously as
@@
When $F\subseteq\overline{C}$ then $C(F)=C$, $I_{C(F)}=I$ and so extending functions by zero gives embeddings of ${\mathbb{X}}_{F}$ and ${\mathbb{X}}_{F}^{\dagger}$ into ${\mathbb{X}}$ which are $P_{F}$-equivariant and $P_{F}^{\dagger}$-equivariant respectively. The algebras ${\mathcal{H}}_{F}$ and ${\mathcal{H}}_{F}^{\dagger}$ can then be naturally identified with subalgebras of ${\mathcal{H}}$ by viewing them as $I$-biinvariant functions on $P_{F}$ and $P_{F}^{\dagger}$ respectively and extending functions by zero. We then have that ${\mathcal{H}}$ is free over ${\mathcal{H}}_{F}^{\dagger}$, which itself is also free over ${\mathcal{H}}_{F}$ (c.f.[25, Lemma 4.20 & Proposition 4.21]).
We also let $G_{\operatorname{aff}}$ be the subgroup of $G$ generated by all the parahoric subgroups $G_{F}$ for $F$ varying among all the faces of $\mathscr{X}$. We can then also construct a $G_{\operatorname{aff}}$-representation ${\mathbb{X}}_{\operatorname{aff}}={\operatorname{ind}}_{I}^{G_{\operatorname{aff}}}(\mathbf{1})$ and an associated Hecke algebra ${\mathcal{H}_{\operatorname{aff}}}={\operatorname{End}}_{G_{\operatorname{aff}}}({\mathbb{X}}_{\operatorname{aff}})$. Via extension by zero, we can once again identify ${\mathbb{X}}_{\operatorname{aff}}$ with a $G_{\operatorname{aff}}$-subrepresentation of ${\mathbb{X}}$ and ${\mathcal{H}_{\operatorname{aff}}}$ with a $k$-subalgebra of ${\mathcal{H}}$. Viewing all the above Hecke algebras as subalgebras of ${\mathcal{H}}$, we note that ${\mathcal{H}_{\operatorname{aff}}}$ contains all the ${\mathcal{H}}_{F}$ for $F\subseteq\overline{C}$ and is in fact generated by them as an algebra (c.f. the proof of [18, Lemma 1.4]). Furthermore, for any $F\subseteq\overline{C}$ we have that ${\mathcal{H}_{\operatorname{aff}}}$ is free as both a left and right ${\mathcal{H}}_{F}$-module.
@@
For any $\tilde{w}\in\widetilde{W}_{M}$, we set $T^{M}_{\tilde{w}}:=I_{M}\tilde{w}I_{M}\in k[I_{M}\backslash M/I_{M}]={\mathcal{H}}_{M}$, the characteristic function of the double coset $I_{M}\tilde{w}I_{M}$. When $M=G$, we drop the superscript and write $T_{\tilde{w}}$. The ring structure of ${\mathcal{H}}_{M}$ is determined by the braid relation
$T^{M}_{\tilde{w}\tilde{w}^{\prime}}=T^{M}_{\tilde{w}}T^{M}_{\tilde{w}^{\prime}}\quad\text{if}\quad\ell_{M}(\tilde{w}\tilde{w}^{\prime})=\ell_{M}(\tilde{w})+\ell_{M}(\tilde{w}^{\prime}),$
and a certain quadratic relation, c.f. [25, §4.8, eq. (4.10)]. By the Bruhat decomposition (2.1), the set $\{T^{M}_{\tilde{w}}\mid\tilde{w}\in\widetilde{W}\}$ is a $k$-basis of ${\mathcal{H}}_{M}$. There are also certain dual elements $T^{M,*}_{\tilde{w}}$ which form another $k$-basis of ${\mathcal{H}}_{M}$, c.f. §4.8, eq. (4.11) in loc. cit. In addition, the subalgebra ${\mathcal{H}_{\operatorname{aff}}}$ also has a $k$-basis given by the set $\{T_{\tilde{w}}\mid\tilde{w}\in\widetilde{W}_{\operatorname{aff}}\}$.
Recall from [1, §4.1] that for $w\in W_{0,M}$ and $\lambda\in\widetilde{\Lambda}$, the element $\lambda\hat{w}\in\widetilde{W}_{M}$ is called $M$-positive (resp. $M$-negative) if $\langle\alpha,\nu(\lambda)\rangle\leq 0$ (resp. $\langle\alpha,\nu(\lambda)\rangle\geq 0$) for every $\alpha\in\Phi^{+}\setminus\Phi_{M}^{+}$. The $k$-span of the $T^{M}_{\tilde{w}}$ with $M$-positive (resp. $M$-negative) $\tilde{w}\in\widetilde{W}$ forms a $k$-subalgebra of ${\mathcal{H}}_{M}$ which we denote by ${\mathcal{H}}_{M}^{+}$ (resp. ${\mathcal{H}}_{M}^{-}$). By Lemma 4.6 of loc. cit. (see also [29, Theorem 1.4(iii)]) we have injective algebra homomorphisms
@@
such that $\iota_{M}({\mathcal{H}}_{M})\cong{\mathcal{H}}_{w_{0}(M)}$.
@@ 2.3. Parabolic induction and simple modules
We recall the construction of the parabolic induction and coinduction functors from [29, §3]. We fix a standard Levi subgroup $M$ of $G$. By (2.4) we may view the algebra ${\mathcal{H}}$ as a bimodule over the algebras ${\mathcal{H}}_{M}^{+}$ and ${\mathcal{H}}_{M}^{-}$ via the maps $j_{M}^{+}$ and $j_{M}^{-}$ respectively. Thus we may define two functors ${\operatorname{Mod}}({\mathcal{H}}_{M})\to{\operatorname{Mod}}({\mathcal{H}})$ by setting
@@
By the tensor-hom adjunction, we see that ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ has left adjoint ${\operatorname{Loc}}_{M}:=(-)\otimes_{{\mathcal{H}}_{M}^{-}}{\mathcal{H}}_{M}$ and ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ has right adjoint ${\mathcal{R}^{G}_{M}}:={\operatorname{Hom}}_{{\mathcal{H}}_{M}^{+}}({\mathcal{H}}_{M},-)$. Note that since ${\mathcal{H}}_{M}$ is the Ore localisation of ${\mathcal{H}}_{M}^{-}$ at $T^{M}_{\lambda_{0}^{-}}$, the functor ${\operatorname{Loc}}_{M}$ is the corresponding localisation functor and hence is exact. The counit of the adjoint pair $({\operatorname{Loc}}_{M},{\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}})$ is given as follows. For ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$, we have a natural map ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})={\operatorname{Hom}}_{{\mathcal{H}}_{M}^{-}}({\mathcal{H}},{\mathfrak{m}})\to{\mathfrak{m}}$ by evaluating at $1$. Localising at $T^{M}_{\lambda_{0}^{-}}$ this gives the counit
@@
[…]
@@
Note that the left (resp. right) adjoint of ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ (resp. ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$) can be described explicitly, namely it is equal to the composite $\mathcal{L}_{M}^{G}:=\iota_{0}^{w_{0}(M)}\circ{\operatorname{Loc}}_{w_{0}(M)}$ (resp. $\iota_{0}^{w_{0}(M)}\circ\mathcal{R}^{G}_{w_{0}(M)}$), c.f. [29, Theorem 1.9].
@@
The next result is, to the best of our knowledge, not explicitly available anywhere in the literature. Its proof was however inspired by the arguments used in the proof of [19, Lemma 8.1]. We note that it also answers the question asked in the last paragraph of [29, §1], in the case where the coefficient ring is a field of characteristic $p$.
@@
For any standard parabolic subgroup $M$ of $G$, the parabolic induction and coinduction functors are fully faithful.
@@
It suffices to show the result for coinduction since they differ by a twist. By 2.4 we must show that the counit $\varepsilon_{{\mathfrak{m}}}$ is an isomorphism for all ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$. By [1, Lemma 4.10] (see also [29, Lemma 3.9]), the map $\varphi\mapsto(\varphi(T_{\hat{v}}))_{v\in W_{0}^{M}}$ gives a $k$-linear isomorphism ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong\bigoplus_{v\in W_{0}^{M}}{\mathfrak{m}}$, where $\hat{v}$ is a fixed lift of $v$ in $\widetilde{W}$ (and by convention we impose $n_{1}=1$). Under this identification and from (2.6), we see that $\varepsilon_{{\mathfrak{m}}}$ is the localisation of the projection onto the $v=1$ summand.
Now, by [1, Proposition 4.12] (whose statement assumes $k$ to be algebraically closed but whose proof works in our level of generality) and [29, Lemma 2.18], the action of $T_{\lambda_{0}^{-}}$ on the summand corresponding to $v\in W_{0}^{M}$ is zero if $\hat{v}^{-1}(\lambda_{0}^{-})$ is not $M$-negative, and the action on the $v=1$ component agrees with the action of $T^{M}_{\lambda_{0}^{-}}$ on ${\mathfrak{m}}$. Hence, from the definition of $\varepsilon_{{\mathfrak{m}}}$, we’ll be done if we show that $\hat{v}^{-1}(\lambda_{0}^{-})$ is not $M$-negative for $v\neq 1$.
@@
We now recall the classification of simple Hecke modules. Here we assume that $k$ is algebraically closed. The classification was initially stated in terms of parabolic coinduction by Abe [1, Theorem 4.22], and was later reworded in terms of parabolic induction (cf. [4, Corollary 4.30]). We first need to recall the definition of a supersingular ${\mathcal{H}}$-module as follows. Recall from [30, §5] that for a choice of spherical orientation $o$, i.e. of a closed Weyl chamber of $W_{0}$, we can associate a basis $\{E_{o}(\tilde{w})\}_{\tilde{w}\in\widetilde{W}}$ of ${\mathcal{H}}$ with the property that $E_{o}(\tilde{w})\in T_{\tilde{w}}+\sum_{\tilde{v}<\tilde{w}}kT_{\tilde{v}}$. Now let $\mathcal{O}$ be a $\widetilde{W}$-conjugacy class in $\tilde{\Lambda}$, and let

### v4

@@ 2.1. The Bruhat-Tits building and Weyl groups
We let $\mathscr{X}$ denote the semisimple Bruhat-Tits building of $G$. In the standard apartment $\mathscr{A}:=X_{*}(\mathbb{T}/\mathbb{C})\otimes_{{\mathbb{Z}}}\mathbb{R}$ of $\mathscr{X}$, we fix a chamber $C$ and a hyperspecial vertex $x_{0}$ such that $x_{0}\in\overline{C}$. Given a face $F$ in $\mathscr{X}$, we let $P_{F}$ denote the parahoric subgroup of $G$ associated to $F$ and $P_{F}^{\dagger}$ the stabiliser in $G$ of the face $F$. The group $P_{F}$ has a pro-$p$ radical which we denote by $I_{F}$. In particular, we set $I=I_{C}$ and $I^{\prime}=P_{C}$ which are called a pro-$p$ Iwahori subgroup and Iwahori subgroup of $G$, respectively.
The chamber $C$ determines a set $\Phi^{+}$ of positive roots of the root system $\Phi\subseteq X^{*}(\mathbb{T}/\mathbb{C})$ of $(\mathbb{G},\mathbb{T})$. We will view elements of $\Phi$ as characters of $T$. If $\Phi^{\prime}$ is an irreducible component of $\Phi$, we write $\mathrm{rk}(\Phi^{\prime})$ to denote its rank. We let $\Pi$ denote the basis of $\Phi$ defined by $\Phi^{+}$, and $B=T\ltimes U$ the Borel subgroup containing $T$ defined by $\Phi^{+}$, where $U$ is the unipotent radical of $B$. A standard parabolic subgroup $P=M\ltimes N$ is any parabolic subgroup of $G$ containing $B$. The corresponding Levi subgroup $M$ will also be called standard. Given a standard parabolic subgroup $P=MN$, we let $\Pi_{M}$ (resp. $\Phi_{M}$, resp. $\Phi_{M}^{+}$) be the corresponding simple roots (resp. root system, resp. positive roots) defined by $M$. We will sometimes also write $\Pi_{P}$ to denote this same set of simple roots.
Recall that the affine functions on $\mathscr{A}$ given by
$(\alpha,m):=\alpha(\bullet)+m\quad\text{for $\alpha\in\Phi$ and $m\in{\mathbb{Z}}$}$
are called the affine roots (note that we are implicitly using splitness of our group, see [27, §4.2] for the general definition). Identifying $\Phi$ with the affine roots of the form $(\alpha,0)$, we will view $\Phi$ as a subset of the set $\Phi_{\operatorname{aff}}$ of affine roots. The chamber $C$ then also determines a set of positive affine roots $\Phi^{+}_{\operatorname{aff}}$ defined as those which take non-negative value on $C$. There is a partial order on $\Phi$ given by $\alpha\leq\beta$ if and only if $\beta-\alpha$ is a non-negative integral linear combination of elements of $\Pi$. We let $\Phi^{\min}\subset\Phi$ denote the set of minimal elements with respect to this partial order, and let $\Pi_{\operatorname{aff}}:=\Pi\cup\{(\alpha,1):\alpha\in\Phi^{\min}\}$.
@@
the extended Weyl group of $(G,T)$, where $\Lambda:=T/(T\cap I^{\prime})$. The action of $G$ on $\mathscr{X}$ restricts to an action of $W$ on $\mathscr{A}$ by affine automorphisms. We let
@@
Let $\Omega=\{w\in W\mid\ell(w)=0\}$. Then we have a decomposition $W=W_{\text{aff}}\rtimes\Omega$ where $W_{\text{aff}}$ denotes the affine Weyl group, generated by the set $S$ of simple affine reflections fixing the walls of $C$. The elements of $S$ are all of the form $s_{\alpha}$ for $\alpha\in\Pi_{\operatorname{aff}}$ (cf. [27, §4.3]). Given any affine root $\alpha=(\beta,m)$ with $\beta\in\Phi$, $m\in{\mathbb{Z}}$, we define $\alpha^{\vee}:=\beta^{\vee}$. Now, given an $\overline{\mathbb{F}_{p}}$-valued character $\xi$ of $T(\mathbb{F}_{q})$, we let
(2.2) $S_{\xi}:=\{s_{\alpha}\in S\mid\xi(\alpha^{\vee}(x))=1\text{ for all }x\in\mathbb{F}_{q}^{\times}\}.$
@@
To each face $F\subseteq\overline{C}$ we associate the subset $S_{F}\subseteq S$ of those reflections fixing $F$ pointwise and we let $W_{F}$ denote the subgroup of $W$ generated by $S_{F}$. Note that $S_{C}=\emptyset$. If $\Phi=\sqcup_{i=1}^{r}\Phi_{i}$ is the decomposition of $\Phi$ into irreducible components and $\Pi=\sqcup_{i=1}^{r}\Pi_{i}$ is the corresponding decomposition of the basis, we also have an associated decomposition $S=\sqcup_{i=1}^{r}S_{i}$. Here $S_{i}=\{s_{\beta}\mid\beta\in\Pi_{i}\}\cup\{s_{(-\alpha_{i},1)}\}$, where $\alpha_{1},\ldots,\alpha_{r}$ denote the highest roots in each irreducible component. Then the map
@@
is a bijection (cf. [8, §1.3.5]). Moreover, this bijection is order-reversing in the sense that $F^{\prime}\subseteq\overline{F}$ if and only if $S_{F}\subseteq S_{F^{\prime}}$.
@@ 2.2. Pro-$p$ Iwahori-Hecke algebras
where $C(F)$ is the chamber defined in [20, Lemma 1.3], and the corresponding Hecke algebras are then defined analogously as
@@
When $F\subseteq\overline{C}$ then $C(F)=C$, $I_{C(F)}=I$ and so extending functions by zero gives embeddings of ${\mathbb{X}}_{F}$ and ${\mathbb{X}}_{F}^{\dagger}$ into ${\mathbb{X}}$ which are $P_{F}$-equivariant and $P_{F}^{\dagger}$-equivariant respectively. The algebras ${\mathcal{H}}_{F}$ and ${\mathcal{H}}_{F}^{\dagger}$ can then be naturally identified with subalgebras of ${\mathcal{H}}$ by viewing them as $I$-biinvariant functions on $P_{F}$ and $P_{F}^{\dagger}$ respectively and extending functions by zero. We then have that ${\mathcal{H}}$ is free over ${\mathcal{H}}_{F}^{\dagger}$, which itself is also free over ${\mathcal{H}}_{F}$ (cf. [27, Lemma 4.20 & Proposition 4.21]).
We also let $G_{\operatorname{aff}}$ be the subgroup of $G$ generated by all the parahoric subgroups $P_{F}$ for $F$ varying among all the faces of $\mathscr{X}$. We can then also construct a $G_{\operatorname{aff}}$-representation ${\mathbb{X}}_{\operatorname{aff}}={\operatorname{ind}}_{I}^{G_{\operatorname{aff}}}(\mathbf{1})$ and an associated Hecke algebra ${\mathcal{H}_{\operatorname{aff}}}={\operatorname{End}}_{G_{\operatorname{aff}}}({\mathbb{X}}_{\operatorname{aff}})$. Via extension by zero, we can once again identify ${\mathbb{X}}_{\operatorname{aff}}$ with a $G_{\operatorname{aff}}$-subrepresentation of ${\mathbb{X}}$ and ${\mathcal{H}_{\operatorname{aff}}}$ with a $k$-subalgebra of ${\mathcal{H}}$. Viewing all the above Hecke algebras as subalgebras of ${\mathcal{H}}$, we note that ${\mathcal{H}_{\operatorname{aff}}}$ contains all the ${\mathcal{H}}_{F}$ for $F\subseteq\overline{C}$ and is in fact generated by them as an algebra (cf. the proof of [20, Lemma 1.4]). Furthermore, for any $F\subseteq\overline{C}$ we have that ${\mathcal{H}_{\operatorname{aff}}}$ is free as both a left and right ${\mathcal{H}}_{F}$-module.
@@
For any $\tilde{w}\in\widetilde{W}_{M}$, we set $T^{M}_{\tilde{w}}:=\mathds{1}_{I_{M}\tilde{w}I_{M}}\in k[I_{M}\backslash M/I_{M}]={\mathcal{H}}_{M}$, the characteristic function of the double coset $I_{M}\tilde{w}I_{M}$. When $M=G$, we drop the superscript and write $T_{\tilde{w}}$. By the Bruhat decomposition (2.1), the set $\{T^{M}_{\tilde{w}}\mid\tilde{w}\in\widetilde{W}\}$ is a $k$-basis of ${\mathcal{H}}_{M}$. There are also certain dual elements $T^{M,*}_{\tilde{w}}$ which form another $k$-basis of ${\mathcal{H}}_{M}$, cf. [27, §4.8, eq. (4.11)]. In addition, the subalgebra ${\mathcal{H}_{\operatorname{aff}}}$ also has a $k$-basis given by the set $\{T_{\tilde{w}}\mid\tilde{w}\in\widetilde{W}_{\operatorname{aff}}\}$. The ring structure of ${\mathcal{H}}$ is determined by the braid relation
$T_{\tilde{w}\tilde{w}^{\prime}}=T_{\tilde{w}}T_{\tilde{w}^{\prime}}\quad\text{if}\quad\ell(\tilde{w}\tilde{w}^{\prime})=\ell(\tilde{w})+\ell(\tilde{w}^{\prime}),$
and a certain quadratic relation which we now describe. Fix the set $\{\hat{s}\mid s\in S\}\subseteq\widetilde{W}_{\operatorname{aff}}$ of lifts of the simple affine reflections defined as in [27, §4.8]. Then the quadratic relations for ${\mathcal{H}}$ are given by
(2.4) $T_{\hat{s}}^{2}=T_{\hat{s}}\cdot\left(|\mu_{\alpha}|\sum_{t\in\alpha^{\vee}(\mathbb{F}_{q}^{\times})}T_{t}\right)$
for all $s=s_{\alpha}\in S$, where $\mu_{\alpha}\cong\ker(\alpha^{\vee}|_{\mathbb{F}_{q}^{\times}})$ has order one or two, cf. [29, §2.2, eqs. (27)-(28), & §2.1.6, eq. (13) & Remark 2.7].
Recall from [1, §4.1] that for $w\in W_{0,M}$ and $\lambda\in\widetilde{\Lambda}$, the element $\lambda\hat{w}\in\widetilde{W}_{M}$ is called $M$-positive (resp. $M$-negative) if $\langle\alpha,\nu(\lambda)\rangle\leq 0$ (resp. $\langle\alpha,\nu(\lambda)\rangle\geq 0$) for every $\alpha\in\Phi^{+}\setminus\Phi_{M}^{+}$. The $k$-span of the $T^{M}_{\tilde{w}}$ with $M$-positive (resp. $M$-negative) $\tilde{w}\in\widetilde{W}$ forms a $k$-subalgebra of ${\mathcal{H}}_{M}$ which we denote by ${\mathcal{H}}_{M}^{+}$ (resp. ${\mathcal{H}}_{M}^{-}$). By Lemma 4.6 of loc. cit. (see also [32, Theorem 1.4(iii)]) we have injective algebra homomorphisms
@@
such that $\iota_{0}^{M}({\mathcal{H}}_{M})\cong{\mathcal{H}}_{w_{0}(M)}$.
@@ 2.3. Parabolic induction and simple modules
We recall the construction of the parabolic induction and coinduction functors from [32, §3]. We fix a standard Levi subgroup $M$ of $G$. By (2.5) we may view the algebra ${\mathcal{H}}$ as an $({\mathcal{H}}_{M}^{+},{\mathcal{H}}_{M}^{-})$-bimodule via the maps $j_{M}^{+}$ and $j_{M}^{-}$ respectively. Thus we may define two functors ${\operatorname{Mod}}({\mathcal{H}}_{M})\to{\operatorname{Mod}}({\mathcal{H}})$ by setting
@@
By the tensor-hom adjunction, we see that ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ has left adjoint ${\operatorname{Loc}}_{M}:=(-)\otimes_{{\mathcal{H}}_{M}^{-}}{\mathcal{H}}_{M}$ and ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ has right adjoint ${\mathcal{R}^{G}_{M}}:={\operatorname{Hom}}_{{\mathcal{H}}_{M}^{+}}({\mathcal{H}}_{M},-)$. Similarly to the above, we are using $j_{M}^{-}$ to view ${\mathcal{H}}$-modules as ${\mathcal{H}}_{M}^{-}$-modules via restriction in the definition of ${\operatorname{Loc}}_{M}$, and similarly $j_{M}^{+}$ is used in the definition of ${\mathcal{R}^{G}_{M}}$. Note as well that since ${\mathcal{H}}_{M}$ is the Ore localisation of ${\mathcal{H}}_{M}^{-}$ at $T^{M}_{\lambda_{0}^{-}}=T^{M,*}_{\lambda_{0}^{-}}$, the functor ${\operatorname{Loc}}_{M}$ is the corresponding localisation functor and hence is exact. The counit of the adjoint pair $({\operatorname{Loc}}_{M},{\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}})$ is given as follows. For ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$, we have a natural map ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})={\operatorname{Hom}}_{{\mathcal{H}}_{M}^{-}}({\mathcal{H}},{\mathfrak{m}})\to{\mathfrak{m}}$ by evaluating at $1$. Localising at $T^{M}_{\lambda_{0}^{-}}$ this gives the counit
@@
    ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}\cong{\operatorname{Coind}}_{{\mathcal{H}}_{w_{0}(M)}}^{{\mathcal{H}}}\circ\iota_{0}^{M}\quad\text{and}\quad{\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}\cong{\operatorname{Ind}}_{{\mathcal{H}}_{w_{0}(M)}}^{{\mathcal{H}}}\circ\iota_{0}^{M},$
@@
Note that the left (resp. right) adjoint of ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ (resp. ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$) can be described explicitly, namely it is equal to the composite $\mathcal{L}_{M}^{G}:=\iota_{0}^{w_{0}(M)}\circ{\operatorname{Loc}}_{w_{0}(M)}$ (resp. $\iota_{0}^{w_{0}(M)}\circ\mathcal{R}^{G}_{w_{0}(M)}$), cf. [32, Theorem 1.9].
@@
The next result is, to the best of our knowledge, not explicitly available anywhere in the literature. Its proof was however inspired by the arguments used in the proof of [21, Lemma 8.1]. We note that it also answers the question asked in the last paragraph of [32, §1], in the case where the coefficient ring is a field of characteristic $p$.
@@
For any standard parabolic subgroup $M$ of $G$, the functors ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ and ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ are fully faithful.
@@
It suffices to show the result for coinduction since the two functors differ by a twist. By 2.4 we must show that the counit $\varepsilon_{{\mathfrak{m}}}$ is an isomorphism for all ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$. By [1, Lemma 4.10] (see also [32, Lemma 3.9]), the map $\varphi\mapsto(\varphi(T_{\hat{v}}))_{v\in W_{0}^{M}}$ gives a $k$-linear isomorphism ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong\bigoplus_{v\in W_{0}^{M}}{\mathfrak{m}}$, where $\hat{v}$ is a fixed lift of $v$ to $\widetilde{W}$ (and by convention we impose for $v=1$ that $\hat{v}=1_{\widetilde{W}}$). Under this identification and from (2.7), we see that $\varepsilon_{{\mathfrak{m}}}$ is the localisation of the projection onto the $v=1$ summand.
Now, we need to consider the action of $T^{*}_{\lambda_{0}^{-}}=j_{M}^{-}(T^{M}_{\lambda_{0}^{-}})$. By [1, Propositions 2.7(ii) & 4.12] (whose statements assume $k$ to be algebraically closed but whose proof works for arbitrary fields of characteristic $p$), the action of $T^{*}_{\lambda_{0}^{-}}$ on the summand corresponding to $v\in W_{0}^{M}$ is zero if $\hat{v}^{-1}(\lambda_{0}^{-})$ is not $M$-negative, and the action on the $v=1$ component agrees with the action of $T^{M}_{\lambda_{0}^{-}}$ on ${\mathfrak{m}}$. Hence, from the definition of $\varepsilon_{{\mathfrak{m}}}$, we’ll be done if we show that $\hat{v}^{-1}(\lambda_{0}^{-})$ is not $M$-negative for $v\neq 1$.
@@
We now recall the classification of simple Hecke modules in terms of supersingulars. The classification was initially stated in terms of parabolic coinduction by Abe [1, Theorem 4.22] for algebraically closed fields, and was later shown to hold for general $k$ in [14, Theorem I.8]. This classification was also reworded in terms of parabolic induction (cf. [4, Corollary 4.30]). We first need to recall the definition of a supersingular ${\mathcal{H}}$-module as follows. Recall from [33, §5] that for a choice of spherical orientation $o$, i.e. of a closed Weyl chamber of $W_{0}$, we can associate a basis $\{E_{o}(\tilde{w})\}_{\tilde{w}\in\widetilde{W}}$ of ${\mathcal{H}}$ with the property that $E_{o}(\tilde{w})\in T_{\tilde{w}}+\sum_{\tilde{v}<\tilde{w}}kT_{\tilde{v}}$. Here $\tilde{v}<\tilde{w}$ refers to the Bruhat order on $W_{\operatorname{aff}}$, suitably extended to $W=W_{\operatorname{aff}}\rtimes\Omega$ and then inflated to $\widetilde{W}$, cf. the discussion preceeding Corollary 5.26 in loc. cit. Now let $\mathcal{O}$ be a $\widetilde{W}$-conjugacy class in $\tilde{\Lambda}$, and let